# She’s Got Her Mother’s Hair: End-to-End Collapsar Simulations Unveil the Origin of Black Holes’ Magnetic Field

### v1

@@ 2.1 Advected flux on the horizon
Once a BH forms, and spins up due to disk accretion, it could conceivably power a relativistic jet capable of escaping the star (MacFadyen & Woosley, 1999). The power of the BZ jet from a Kerr BH of mass $\,{M_{\rm BH}}$ depends on both its (dimensionless) spin $\,{a_{\rm BH}}$ and the magnetic flux according to (Tchekhovskoy et al., 2011; Gottlieb et al., 2023b):
@@
where $\eta_{a}=1.063\,{a_{\rm BH}}^{4}+0.395\,{a_{\rm BH}}^{2}$ (Lowell et al., 2024) and $\eta_{\phi}=(\Phi/\phi_{\rm MAD})^{2}(\dot{M}\,{r_{\rm g}}^{2}c)^{-1}$ are the spin and flux efficiencies, respectively, and $\,{r_{\rm g}}$ is the BH gravitational radius. Here, $\phi_{\rm MAD}\approx 50$ is the maximum dimensionless magnetic flux (Tchekhovskoy et al., 2011), corresponding to a magnetically-arrested disk (MAD). Thus, even for a rapidly spinning BH, a magnetic flux $\Phi\gtrsim 10^{27}$ G cm2 is required to explain the typical luminosities $L_{\rm j}\gtrsim 10^{49}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ of GRB jets (e.g., Butler et al. 2010; Wanderman & Piran 2010). As illustrated by the red shaded region in the bottom panel of Fig. 2, which assumes $\,{a_{\rm BH}}=0.5$ and $M_{\mathrm{BH}}=2.7\,\,{M_{\odot}}$, only the most unrealistically optimistic scenario in our progenitor model can provide such a flux. In the realistic scenario of canceling polarities, the required magnetic flux falls short by several orders of magnitude. We conclude that the magnetic fields generated in the progenitor star are likely insufficient for the BH to launch a relativistic jet.
@@ 3.1 Preserving the PNS’s magnetic field
Having established the progenitor star to be an unlikely source of magnetic flux, we now consider the conditions under which the nascent BH can inherit the magnetic field of the PNS. For this to occur, the PNS magnetosphere must be anchored close to the PNS surface by an accretion disk prior to and following BH formation. Absent such a disk, the newly formed BH will quickly shed its magnetic field. Resistive GRMHD simulations find this “balding” process to occur on a timescale (e.g., Bransgrove et al., 2021; Selvi et al., 2024):
@@ 3.2 Natal black hole spin and magnetic field
The maximum magnetic flux a BH can acquire from a PNS reflects that of the PNS prior to collapse. In the presence of an $\alpha-\omega$ dynamo (e.g., Thompson & Duncan 1993), the large-scale magnetic field of the PNS depends on a comparison between its convective overturn time $\tau_{c}$ and rotational period, $P$. For rapidly spinning progenitor stars of interest, the latter can range from $P\sim\,{{\rm ms}}$ close to the NS surface to much higher values $P\gg 1\,\,{{\rm ms}}$ at greater depths, corresponding to a broad range in the Rossby number throughout the PNS. Raynaud et al. (2020); Aloy & Obergaulinger (2021); Barrère et al. (2024) found $Ro\lesssim 1$ close to the NS surface, supporting an efficient convective dynamo capable of PNS large-scale field amplification by orders of magnitude to $B\approx 10^{15}$ G.
The spin of the BHs created in the majority of (failed) core-collapse supernovae is generally expected to be moderate. Specifically, spin values of $0.2\lesssim\,{a_{\rm BH}}\lesssim 0.5$ are needed to power GRB jets through the BZ mechanism (Gottlieb et al., 2023a). A PNS rotating at period $P$ at the time of collapse will form a BH of spin:
@@
where $M$ and $I=0.35M\,{{R_{\rm NS}}}^{2}$ are the PNS mass and moment of inertia, the latter is calculated following Bejger & Haensel (2002) for an assumed compactness $0.12$ (Dietrich et al., 2020).
After forming but prior to collapse, the PNS rotation rate evolves due to a combination of accretion and magnetic braking. Insofar that the latter only becomes relevant once the magnetic field is amplified, three distinct timescales enter the problem: the PNS collapse time $\,{{t_{\rm BH}}}$, the magnetic field amplification timescale $\,{{t_{\rm B}}}$, and the magnetic braking spin-down time, $\,{{t_{\rm spin}}}$.
@@
where the factor $(1-\Omega/\Omega_{k})$ limits the PNS to spin below centrifugal break-up velocity, $\Omega_{k}$ is now the Keplerian angular velocity at the PNS surface, and
@@
where we have set $M=\,{M_{\rm BH}}$. Eq. (3.2) implies that the maximum equilibrium spin, as obtained by the PNS break-up velocity, is $a_{\rm BH,max}\approx 0.58\sqrt{R_{12}/M_{3}}$, where $R_{12}=R/12\,\,{{\rm km}}$ and $M_{3}=M/3\,\,{M_{\odot}}$. This result is consistent with full numerical solutions of rotating hydrostatic equilibrium (Margalit et al., 2015).
Figure 3: Mapping of the PNS poloidal magnetic field strength (left axis), flux (right axis), and [assumed temporally constant] mass accretion rate onto the PNS (horizontal axis) to the natal BH spin (color map), PNS spin-down energy released prior to BH formation (yellow contours, in logarithmic units) and logarithmic BZ jet power (black contours), assuming initial $P_{0}=1\,\,{{\rm ms}}$. The BH cannot inherit the PNS field if the Alfvén radius fails to satisfy $R_{A}\lesssim\,{{R_{\rm NS}}}$ (the excluded area is in gray). The dotted vertical (dashed horizontal) black lines delineate the estimated range of mass accretion rates (PNS dipole field) at the time of collapse from PNS simulations by Aloy & Obergaulinger (2021) (Raynaud et al., 2020). The moderate spin leads to jets with a characteristic GRB power. The MAD state with $\eta_{\phi}\approx 1$ is located deep in the gray zone where the BH does not acquire the PNS field. The natal BH spin distribution does not change significantly for $P_{0}=10\,\,{{\rm ms}}$.
@@
During the PNS spin-down phase, PNS winds are ejected into the collapsing star, inhibiting accretion along the polar axis and potentially preventing the magnetic field lines from being drawn into the BH upon collapse. The yellow contours delineate the logarithm of the spin-down energy emitted by the PNS before the collapse to a BH. Given the estimated accretion rate and dipole magnetic field of the PNS, the spin-down energy is $\sim 10^{51}-10^{52}\,\,{{\rm erg}}$, aligns with the excess energy observed in SNe Ic-BL associated with collapsar GRBs (e.g., Cano et al., 2017).
@@
In the opposite regime of $\,{{t_{\rm BH}}}\lesssim\,{{t_{\rm B}}}$, there might not be enough time for the field to reach saturation before the PNS collapses. In such cases, the BH may form before an appreciable spin-down of the PNS. As a result, the BH will likely acquire a dimensionless spin similar to $P_{0}$. We conclude that for expected $1\,\,{{\rm ms}}\lesssim P_{0}\lesssim 10\,\,{{\rm ms}}$, the natal BH spin is moderate, irrespective of $\,{{t_{\rm B}}}/\,{{t_{\rm BH}}}$, and the acquired field is less than the saturation level but may still reach high values depending on $\,{{t_{\rm B}}}/\,{{t_{\rm BH}}}$. If $\,{{t_{\rm B}}}\lesssim\,{{t_{\rm BH}}}$, the nascent BH forms with sufficient magnetic flux to launch a GRB-like jet immediately.
@@ 4.1 GRMHD setup
We conduct a suite of GRMHD simulations utilizing the 3D GPU-accelerated code h-amr (Liska et al., 2022), leveraging an ideal equation of state featuring an adiabatic index of $4/3$. This equation of state is well-suited for modeling the radiation-dominated gas prevalent in the core. The simulations consider BHs embedded in a stellar envelope to investigate the evolution of the remnant magnetic field on the BH as a function of the BH spin, magnetic field, and the presence of a disk.
In §3 established that the BH likely forms with a moderate spin. However, depending on the value of $\eta_{\phi}$, it may either spin down or spin up at later times. Consequently, we vary the dimensionless spin parameter across different simulation models. We emphasize that our simulations assume a static metric, maintaining a fixed BH spin throughout. A non-negligible spin-down within the simulation time $T_{s}\gtrsim 6\,\,{{\rm s}}$ is expected if the initial BH spin is close to unity (Jacquemin-Ide et al., 2024). This implies that the simulated jets powered by rapidly spinning BHs will be somewhat stronger than if the spin-down effect were accounted for.
As our simulations consistently feature a central BH engine, we assume that the magnetic field of the collapsing PNS threads the BH upon formation, as outlined in §3. To initiate the BH with a strong magnetic field, we initialize the BH vicinity with a vertical magnetic field and total magnetic flux equivalent to that of the PNS. This approach facilitates early accumulation of the magnetic flux on the BH, with subsequent field evolution governed by the magnetohydrodynamics in the BH vicinity. Specifically, at the onset of the simulation, the core hosts a constant poloidal field of $B_{0}=10^{12.5}\,{\rm G}$. The radial extent of the core magnetic field, determined by flux freezing, is given by $R_{c}=\,{{R_{\rm NS}}}\left(B_{\rm dip}/B_{0}\right)^{1/2}$. Despite initializing the flux with $\Phi_{0}\gtrsim 10^{28}\,{\rm G\,\,{{\rm cm}}^{2}}$, stochastic processes lead to a reduced flux on the BH, $\Phi_{s}\lesssim 0.3\Phi_{0}$, depending on the specific setup. Table 1 outlines the different simulations considered here.
@@
We set $\,{M_{\rm BH}}=2.7\,\,{M_{\odot}}$ as an estimate of the maximum PNS mass, which varies with rotation and equation of state (see discussion in Margalit & Metzger, 2017). To model the spherically-symmetric density profile in our simulations, we semi-analytically evolve the gas free-fall of the progenitor presented in §2 from the onset of core-collapse to the BH formation time, $\sim 1\,\,{{\rm s}}$ post-core-collapse (see vertical dashed black line in Fig. 2). Hence, at the beginning of our simulations, the density profiles adhere to free-fall profiles with a power-law index of $-1.5$. We neglect the gas thermal pressure to compensate for the absence of self-gravity in the simulations.
For the grid setup, we employ spherical polar coordinates, $r$, $\theta$, $\phi$. The $r$-direction is logarithmically distributed, from slightly inside the event horizon to $4\times 10^{4}\,\,{{\rm km}}$. Uniform cell size distribution is maintained along the $\theta$- and $\phi$-directions. The basic grid encompasses $N_{r}\times N_{\theta}\times N_{\phi}$ cells in the $r$-, $\theta$-, and $\phi$-directions, respectively, where $N_{r}=256$; $N_{\theta}=128$; and $N_{\phi}=128$. We incorporate local adaptive time-stepping and static mesh refinement to enhance computational efficiency. For all simulations, refinement is implemented at one level within the innermost $r=80\,\,{r_{\rm g}}$ to ensure proper resolution of the wavelength of the fastest-growing MRI mode.
@@ 4.2 Results
As shown in §2, the accretion disk is anticipated to be present at the time of PNS collapse. Therefore, we simulate the collapse of a star with a rotation profile that facilitates the rapid formation of an accretion disk, before the total magnetic flux accumulates on the BH. In simulations with higher (lower) magnetic flux, most of the initial flux is advected onto the BH within $\sim 0.2~{}(0.1)\,\,{{\rm s}}$. Over the following $\sim 0.1\,\,{{\rm s}}$, some magnetic flux is lost to reconnection and stabilizes with $\Phi_{s}\approx 0.3\Phi_{0}$ threading the BH, if a disk is present. If there is no disk to hamper reconnection, the flux drops to $\Phi_{s}\lesssim 0.1\Phi_{0}$.
@@
A relativistic jet is also observed in the models in the second row, where either a rapidly spinning BH and a lower initial flux ($a9\Phi lD$) or a slowly spinning BH with a higher initial flux ($a1\Phi hD$). However, in these configurations, the jet exhibits diminished strength attributed to the reduced field intensity or the slower rotation of the BH, compared to model $a9\Phi hD$. This is demonstrated by the field line density along the relativistic jets in Fig. 4. As a result, the jet manifests intermittency, with instances at which no jet is observed emanating from the BH, or only a one-sided jet is launched.
@@
Figure 5 illustrates the temporal evolution of various parameters on the horizon. The BH maintains a strong magnetic field (a) and high aggregated flux (b) so long as an accretion disk is present, irrespective of the BH spin. However, models with low spin and low flux ($a1\Phi lD$) or without a disk struggle to sustain strong fields over long times. In Fig. 5(c), which depicts the flux on the northern hemisphere, it is apparent that models with weak fields have their magnetic flux reconnects, forming opposite polarity loops threading through the horizon (dashed lines depict negative polarity). Even in the absence of the disk, some magnetic flux remains due to the quasi-spherical accretion that retains it, although it lacks a distinct dipolar field to launch jets.
The dimensionless magnetic flux (d) reveals that the flux is preserved in models with initially high flux and where an accretion disk is present. None of the models exhibit a MAD state ($\phi<\phi_{\rm MAD}$), which is obtained only for very low mass accretion rates and strong magnetic fields (Fig. 3). However, a MAD state might be achieved as the mass accretion rate will ultimately go down. The spin efficiency plays a crucial role in the total electromagnetic launching efficiency (e) and jet luminosity (f). When both the flux and the spin are high ($a9\Phi hD$), the jet power surpasses that of the observed GRB population (see Gottlieb et al., 2023a). Therefore, as suggested in §3, the BHs that power GRB jets should have either a lower magnetic flux or spin, $\,{a_{\rm BH}}\lesssim 0.5$. Indeed, models featuring initially lower magnetic field strengths or moderate BH spins yield jets with power consistent with observations. Conversely, insufficient BH spin or initial field strength, or the absence of a confining disk, fails to produce relativistic outflows.

### v2

@@ 2.1 Advected flux on the horizon
Once a BH forms and spins up due to disk accretion, it could conceivably power a relativistic jet capable of escaping the star (MacFadyen & Woosley, 1999). The power of the BZ jet from a Kerr BH of mass $\,{M_{\rm BH}}$ depends on both its (dimensionless) spin $\,{a_{\rm BH}}$ and the magnetic flux according to (Tchekhovskoy et al., 2011; Gottlieb et al., 2023b):
@@
where $\eta_{a}=1.063\,{a_{\rm BH}}^{4}+0.395\,{a_{\rm BH}}^{2}$ (Lowell et al., 2024) and $\eta_{\phi}=(\Phi/\phi_{\rm MAD})^{2}(\dot{M}\,{r_{\rm g}}^{2}c)^{-1}$ are the spin and flux efficiencies, respectively, and $\,{r_{\rm g}}$ is the BH gravitational radius. Here, $\phi_{\rm MAD}\approx 50$ is the maximum dimensionless magnetic flux (Tchekhovskoy et al., 2011), corresponding to a magnetically arrested disk (MAD). Thus, even for a rapidly spinning BH, a magnetic flux $\Phi\gtrsim 10^{27}$ G cm2 is required to explain the typical luminosities $L_{\rm j}\gtrsim 10^{49}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ of GRB jets (e.g., Butler et al. 2010; Wanderman & Piran 2010). As illustrated by the red shaded region in the bottom panel of Fig. 2, which assumes $\,{a_{\rm BH}}=0.5$ and $M_{\mathrm{BH}}=2.7\,\,{M_{\odot}}$, only the most unrealistically optimistic scenario in our progenitor model can provide such a flux. In the realistic scenario of canceling polarities, the required magnetic flux falls short by several orders of magnitude. We conclude that the magnetic fields generated in the progenitor star are likely insufficient for the BH to launch a relativistic jet.
@@ 3.1 Preserving the PNS’s magnetic field
Having established the progenitor star to be an unlikely source of magnetic flux, we now consider the conditions under which the nascent BH can inherit the magnetic field of the PNS. For this to occur, the PNS magnetosphere must be anchored close to the PNS surface by an accretion disk prior to and following BH formation. Absent such a disk, the newly formed BH will quickly shed its magnetic field. Resistive GRMHD simulations find that this “balding” process occurs on a timescale (e.g., Bransgrove et al., 2021; Selvi et al., 2024):
@@ 3.2 Natal black hole spin and magnetic field
The maximum magnetic flux a BH can acquire from a PNS reflects that of the PNS prior to collapse. In the presence of an $\alpha-\omega$ dynamo (e.g., Thompson & Duncan 1993), the large-scale magnetic field of the PNS depends on a comparison between its convective overturn time $\tau_{c}$ and rotational period, $P$. For rapidly spinning progenitor stars of interest, the latter can range from $P\sim\,{{\rm ms}}$ close to the NS surface to much higher values $P\gg 1\,\,{{\rm ms}}$ at greater depths, corresponding to a broad range in the Rossby number throughout the PNS. Raynaud et al. (2020); Aloy & Obergaulinger (2021) found $Ro\lesssim 1$ close to the NS surface, supporting an efficient convective dynamo capable of PNS large-scale field amplification by orders of magnitude to $B\approx 10^{15}$ G.
The spin of the BHs created in the majority of (failed) core-collapse SNe is generally expected to be moderate. Specifically, spin values of $0.2\lesssim\,{a_{\rm BH}}\lesssim 0.5$ are needed to power GRB jets through the BZ mechanism (Gottlieb et al., 2023a). A PNS rotating at period $P$ at the time of collapse will form a BH of spin:
@@
where $M$ and $I=0.35M\,{{R_{\rm NS}}}^{2}$ are the PNS mass and moment of inertia, the latter is calculated following Bejger & Haensel (2002) for an assumed compactness of $0.12$ (Dietrich et al., 2020).
After forming but prior to collapse, the PNS rotation rate evolves due to a combination of accretion and magnetic braking. Insofar as the latter only becomes relevant once the magnetic field is amplified, three distinct timescales enter the problem: the PNS collapse time $\,{{t_{\rm BH}}}$, the magnetic field amplification timescale $\,{{t_{\rm B}}}$, and the magnetic braking spin-down time, $\,{{t_{\rm spin}}}$.
@@
where the factor $(1-\Omega/\Omega_{k})$ limits the PNS to spin below the centrifugal breakup velocity, $\Omega_{k}$ is now the Keplerian angular velocity at the PNS surface, and
@@
where we have set $M=\,{M_{\rm BH}}$. Eq. (3.2) implies that the maximum equilibrium spin, as obtained by the PNS breakup velocity, is $a_{\rm BH,max}\approx 0.58\sqrt{R_{12}/M_{3}}$, where $R_{12}=R/12\,\,{{\rm km}}$ and $M_{3}=M/3\,\,{M_{\odot}}$. This result is consistent with full numerical solutions of rotating hydrostatic equilibrium (Margalit et al., 2015).
Figure 3: Mapping of the PNS poloidal magnetic field strength (left axis), flux (right axis), and [assumed temporally constant] mass accretion rate onto the PNS (horizontal axis) to the natal BH spin (color map), PNS spin-down energy released prior to BH formation (yellow contours; in logarithmic units) and logarithmic BZ jet power (black contours), assuming initial $P_{0}=1\,\,{{\rm ms}}$. The BH cannot inherit the PNS field if the Alfvén radius fails to satisfy $R_{A}\lesssim\,{{R_{\rm NS}}}$ (the excluded area is in gray). The dotted vertical (dashed horizontal) black lines delineate the estimated range of mass accretion rates (PNS dipole field) at the time of collapse from PNS simulations by Aloy & Obergaulinger (2021) (Raynaud et al., 2020). The moderate spin leads to jets with a characteristic GRB power. The MAD state with $\eta_{\phi}\approx 1$ is located deep in the gray zone, where the BH does not acquire the PNS field. The natal BH spin distribution does not change significantly for $P_{0}=10\,\,{{\rm ms}}$.
@@
During the PNS spin-down phase, PNS winds are ejected into the collapsing star, inhibiting accretion along the polar axis and potentially preventing the magnetic field lines from being drawn into the BH upon collapse. The yellow contours delineate the logarithm of the spin-down energy emitted by the PNS before the collapse to a BH. Given the estimated accretion rate and dipole magnetic field of the PNS, the spin-down energy is $\sim 10^{51}-10^{52}\,\,{{\rm erg}}$, and aligns with the excess energy observed in Type Ic-BL SNe associated with collapsar GRBs (e.g., Cano et al., 2017).
@@
In the opposite regime of $\,{{t_{\rm BH}}}\lesssim\,{{t_{\rm B}}}$, there might not be enough time for the field to reach saturation before the PNS collapses. In such cases, the BH may form before an appreciable spin-down of the PNS. As a result, the BH will likely acquire a dimensionless spin similar to $P_{0}$. We conclude that for the expected $1\,\,{{\rm ms}}\lesssim P_{0}\lesssim 10\,\,{{\rm ms}}$, the natal BH spin is moderate, irrespective of $\,{{t_{\rm B}}}/\,{{t_{\rm BH}}}$, and the acquired field is less than the saturation level but may still reach high values depending on $\,{{t_{\rm B}}}/\,{{t_{\rm BH}}}$. If $\,{{t_{\rm B}}}\lesssim\,{{t_{\rm BH}}}$, the nascent BH forms with sufficient magnetic flux to launch a GRB-like jet immediately.
@@ 4.1 GRMHD setup
We conduct a suite of GRMHD simulations utilizing the 3D GPU-accelerated code h-amr (Liska et al., 2022), leveraging an ideal equation of state featuring an adiabatic index of $4/3$. This equation of state is well-suited for modeling the radiation-dominated gas prevalent in the core. The simulations consider BHs embedded in a stellar envelope to investigate the evolution of the remnant magnetic field on the BH as a function of the BH spin, magnetic field, and presence of a disk.
In §3, we established that the BH likely forms with a moderate spin. However, depending on the value of $\eta_{\phi}$, it may either spin down or spin up at later times. Consequently, we vary the dimensionless spin parameter across different simulation models. We emphasize that our simulations assume a static metric, maintaining a fixed BH spin throughout. A non-negligible spin-down within the simulation time $T_{s}\gtrsim 6\,\,{{\rm s}}$ is expected if the initial BH spin is close to unity (Jacquemin-Ide et al., 2024). This implies that the simulated jets powered by rapidly spinning BHs will be somewhat stronger than if the spin-down effect were accounted for.
As our simulations do not model the PNS phase and consistently feature a central BH engine, we assume that the magnetic field of the collapsing PNS threads the BH upon formation, as outlined in §3. To initiate the BH with a strong magnetic field, we initialize the BH vicinity with a vertical magnetic field and total magnetic flux equivalent to that of the PNS. This approach facilitates early accumulation of the magnetic flux on the BH, with subsequent field evolution governed by the magnetohydrodynamics in the BH vicinity. Specifically, at the onset of the simulation, the core hosts a constant poloidal field of $B_{0}=10^{12.5}\,{\rm G}$. The radial extent of the core magnetic field, determined by flux freezing, is given by $R_{c}=\,{{R_{\rm NS}}}\left(B_{\rm dip}/B_{0}\right)^{1/2}$. Despite initializing the flux with $\Phi_{0}\gtrsim 10^{28}\,{\rm G\,\,{{\rm cm}}^{2}}$, stochastic processes lead to a reduced flux on the BH, $\Phi_{s}\lesssim 0.3\Phi_{0}$, depending on the specific setup. Table 1 outlines the different simulations considered here.
@@
We set $\,{M_{\rm BH}}=2.7\,\,{M_{\odot}}$ as an estimate of the maximum PNS mass, which varies with rotation and equation of state (see discussion in Margalit & Metzger, 2017). Our simulations do not include the core collapse and the PNS phase, which will affect the gas distribution in the collapsing star. Therefore, we initialize a spherically symmetric density profile by semi-analytically evolving the gas free-fall of the progenitor presented in §2 from the onset of core collapse to the BH formation time, $\sim 1\,\,{{\rm s}}$ post-core collapse (see vertical dashed black line in Fig. 2). Hence, at the beginning of our simulations, the density profiles adhere to free-fall profiles with a power-law index of $-1.5$. We neglect the gas thermal pressure to compensate for the absence of self-gravity in the simulations.
For the grid setup, we employ spherical polar coordinates, $r$, $\theta$, $\varphi$. The $r$-direction is logarithmically distributed, from slightly inside the event horizon to $4\times 10^{4}\,\,{{\rm km}}$. Uniform cell size distribution is maintained along the $\theta$- and $\varphi$-directions. The basic grid encompasses $N_{r}\times N_{\theta}\times N_{\varphi}$ cells in the $r$-, $\theta$-, and $\varphi$-directions, respectively, where $N_{r}=256$; $N_{\theta}=128$; and $N_{\varphi}=128$. We incorporate local adaptive time-stepping and static mesh refinement to enhance computational efficiency. For all simulations, refinement is implemented at one level within the innermost $r=80\,\,{r_{\rm g}}$ to ensure proper resolution of the wavelength of the fastest-growing MRI mode.
@@ 4.2 Results
As shown in §2, the excess angular momentum in the core implies that the accretion disk will be present during the PNS collapse. Therefore, we simulate the collapse of a star with a rotation profile that facilitates the rapid formation of an accretion disk, before the total magnetic flux accumulates on the BH. In simulations with higher (lower) magnetic flux, most of the initial flux is advected onto the BH within $\sim 0.2~{}(0.1)\,\,{{\rm s}}$. Over the following $\sim 0.1\,\,{{\rm s}}$, some magnetic flux is lost to reconnection and stabilizes with $\Phi_{s}\approx 0.3\Phi_{0}$ threading the BH, if a disk is present. If there is no disk to hamper reconnection, the flux drops to $\Phi_{s}\lesssim 0.1\Phi_{0}$.
@@
A relativistic jet is also observed in the models in the second row, where there is either a rapidly spinning BH with a lower initial flux ($a9\Phi lD$) or a slowly spinning BH with a higher initial flux ($a1\Phi hD$). However, in these configurations, the jet exhibits diminished strength attributed to the reduced field intensity or the slower rotation of the BH, compared to model $a9\Phi hD$. This is demonstrated by the field line density along the relativistic jets in Fig. 4. As a result, the jet manifests intermittency, with instances in which no jet is observed emanating from the BH, or only a one-sided jet is launched.
@@
Figure 5 illustrates the temporal evolution of various parameters on the horizon. The BH maintains a strong magnetic field (a) and high aggregated flux (b) so long as an accretion disk is present, irrespective of the BH spin. However, models with low spin and low flux ($a1\Phi lD$) or without a disk struggle to sustain strong fields over long times. In Fig. 5(c), which depicts the flux on the northern hemisphere, it is apparent that models with weak fields have their magnetic flux reconnect, forming opposite polarity loops threading through the horizon (dashed lines depict negative polarity). Even in the absence of the disk, some magnetic flux remains due to the quasi-spherical accretion that retains it, although it lacks a distinct dipolar field to launch jets.
The dimensionless magnetic flux (d) reveals that the flux is preserved in models with initially high flux and where an accretion disk is present. None of the models exhibit a MAD state ($\phi<\phi_{\rm MAD}$), which is obtained only for very low mass accretion rates and strong magnetic fields (Fig. 3). However, a MAD state might be achieved as the mass accretion rate will ultimately go down. The spin efficiency plays a crucial role in the total electromagnetic launching efficiency (e) and jet luminosity (f). When both the flux and the spin are high ($a9\Phi hD$), the jet power surpasses that of the observed GRB population (see Gottlieb et al., 2023a). Therefore, as suggested in §3, the BHs that power GRB jets should have either a lower magnetic flux or spin, $\,{a_{\rm BH}}\lesssim 0.5$. Indeed, models featuring initially lower magnetic field strengths or moderate BH spins yield jets with power consistent with observations. Conversely, insufficient BH spin or initial field strength or the absence of a confining disk fail to produce relativistic outflows.